OpenStudy (anonymous):

complex no.,,, what the arg z of 0+1i??

OpenStudy (nikvist):

\[\arg(z)=\pi/2\]

OpenStudy (anonymous):

can u prove it for me??

OpenStudy (nikvist):

\[e^{i\phi}=\cos{\phi}+i\sin{\phi}=i\quad\Rightarrow\quad\cos{\phi}=0\quad,\quad\sin{\phi}=1\quad\Rightarrow\quad\phi=\pi/2\]

OpenStudy (anonymous):

tq,